# Optimize Weight Rounding via Signed Gradient Descent for the Quantization of LLMs

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional proficiency in language-related tasks. However, their deployment presents significant challenges due to their substantial memory and storage requirements. To address this challenge, weight-only quantization has emerged as a promising solution. Previous research has indicated that fine-tuning through up and down rounding can enhance performance. In this study, we introduce SignRound, a method that utilizes signed gradient descent (SignSGD) to optimize rounding values and weight clipping within just 200 steps, combining the strengths of both Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). SignRound achieves outstanding results compared to recent methods across 2 to 4 bits, while maintaining low tuning costs and without introducing any additional inference overhead. For instance, SignRound led to absolute average accuracy improvements ranging from 6.91% to 33.22% at 2 bits. Furthermore, it demonstrates robust generalization to various recent models and achieves near-lossless quantization in most scenarios at 4 bits. The source code is publicly available at https://github.com/intel/auto-round.
@@ 4 Experiments
In this section, we conduct a comprehensive evaluation of SignRound from various perspectives. Initially, we offer a concise overview of the LLM architectures and tasks encompassed in our assessment. Next, we delve into a detailed comparison between our method and several existing approaches, highlighting the unique advantages of SignRound. Furthermore, we conduct ablation studies to strengthen the efficacy of our choices and explore the sensitivity of hyperparameters. Lastly, we assess the generation ability of our method across various recent models. The tuning cost comparisons are present in Appendix A.
@@ 4.1 Experimental Settings
We evaluate multiple language tasks to accommodate the task-agnostic setting. Specifically, we present the average accuracy results for 11 zero shot tasks, including HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), LAMBADA (Paperno et al., 2016), TruthfulQA (Lin et al., 2021), OpenBookQA (Mihaylov et al., 2018), BoolQ (Clark et al., 2019), RTE (Dagan et al., 2010), ARC-Easy, ARC-Challenge (Clark et al., 2018) and MMLU (Hendrycks et al., 2020). We use lm-eval-harness(Gao et al., 2023) for all the above tasks. Furthermore, we complement our evaluation with perplexity (ppl) analysis on Wikitext2 (Merity et al., 2016) PTB (Marcus et al., 1993) and C4 (Raffel et al., 2020), by following the source code  [^1] https://github.com/IST-DASLab/gptq of GPTQ and Wikitext2 (Merity et al., 2016) using lm-eval-harness(Gao et al., 2023). However, We argue that perplexity is notably influenced by outliers,as illustrated in Table 14 for different algorithms. This susceptibility likely arises from the mathematical expression $\mathrm{PPL}(X)=\exp\left(-\frac{1}{t}\sum_{i=1}^{t}\log p_{\theta}(x_{i}|x_{<i})\right)$, where assigning a low probability to even one token can significantly inflate the perplexity score. Consequently, we prioritize the accuracy of the 11 tasks mentioned above as the primary metric, with perplexity data serving as a supplementary reference.
@@
In accordance with GPTQ (Frantar et al., 2022), our focus lies specifically on weight-only quantization, targeting the linear layers within transformer blocks. Layers such as the embedding layer and typically the last linear layer like ’lm-head’ are excluded from the quantization process. Our evaluation primarily centers on W4, W4G128, and W3G128 configurations, where W4 indicates quantizing weights with 4 bits and G represents finer-grained grouping as described in (Park et al., 2022; Frantar et al., 2022). We adopt asymmetric quantization for these configurations. For calibration and to mitigate overfitting on the WikeText and C4 datasets, we randomly select 512 samples with the same seed from the readily available pile-10k dataset  [^2] https://huggingface.co/datasets/NeelNanda/pile-10k, which comprises the first 10k samples from pile (Gao et al., 2020). We used a sequence length of 2048 for calibration, while for other methods, we adhere to their official settings
@@
We compare different algorithms on LLaMA-V1 (Touvron et al., 2023a), LLaMA-V2 (Touvron et al., 2023b) and Mistral-7B-v0.1 (Jiang et al., 2023) as they are commonly used. We cover a wide range of LLM parameters, ranging from 7B to 70B, to ensure comprehensive coverage and analysis.
Table 1: Average accuracies($\uparrow$) across 11 tasks, detailed in Section 4.1, of LLaMA and Mistral models at W2-W4. Our* denotes the highest accuracy achieved among the 8 hyperparameter choices, detailed in Section 4.2 , whereas for the 70B model, we only tested a few options.
@@
Table 2: Detrailed accuracies($\uparrow$) across 11 tasks(0-shot) of LLaMA and Mistral models at W4G-1. Our* denotes the highest accuracy achieved among the 8 hyperparameter choices, whereas for the 70B model, we only tested a few options, detailed in 4.2. Appendix C provides more detailed data.
@@
SignRound Hyper-parameters.
Unless explicated stated, the tuning process involved adjusting each block for 200 steps , with the learning rate(lr) as 1/steps, i.e. $35e-$ 3, a batch size of 8 and linear lr decay. Additionally, we employed automatic mixed precision (AMP) to accelerate the tuning.
@@ 4.2 Comparing With Recent Methods
In this section, we compare our methods with those that have already demonstrated remarkable results and impose no additional overhead on our tested models in weight-only quantization for LLMs, including GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2023), HQQ (Badri and Shaji, 2023), Omniquant, (Shao et al., 2023) with a naive method RTN.
To ensure fair comparison as much as possible, we have enabled act-order and true-sequential in GPTQ, and also activated static_group in scenarios where $group\_size\neq-1$. The notation GPTQ+ indicates that we adjusted the random seed or data pre-processing to address issues related to the non-positive definite Hessian matrix or other issues. For Omniquant(Shao et al., 2023), we adhere to the official settings, which include running for 20 epochs including W2G128 for saving time and disabling ’let’. We conducted calibration tests using sample sizes of 512 and 128, as well as a sample size of 512 with a batch size of 4. Our findings show that using a sample size of 512 typically results in comparable or slight higher performance for models $<=13B$. Therefore, we present the results based on the sample size of 512. For 70B models, due the the NAN loss issue and to reduce the tuning cost of Omniquant, we adopted 128 samples for calibration.
We present the summary results of Mistral-7B and LLAMA V2 in Table 1 , detailed results of W4G-1 in Table 2 Additional detailed results are provided in Appendix C due to space constraints. In summary, our approach demonstrated superior performance compared to GPTQ (Frantar et al., 2022), achieving scores of 30/32, AWQ (Lin et al., 2023) with 27/32, HQQ (Badri and Shaji, 2023) with 15/16, and OmniQuant (Shao et al., 2023) with a score of 29/32 across llamv1/llamav2/mistral-7b on various quantization settings, including W4G-1, W4G128, W3G128, and W2G128. These evaluations were based on the average accuracies of 11 zero-shot tasks.
It’s worthy to noting that, as the bits going down, the advantages of Signround becomes more notable. For example, at W2G128, SignRound could yield absolute average accuracy improvements ranging from 6.91% to 33.22%.
Moreover, we can enhance the performance by optimizing the model’s hyperparameters from a selection of eight choices, denoted as ours*. These include steps (200, 1000), weight clip lr (1.0/steps, 2.0/steps), and the option to either enable or disable quanted inputs which means utilizing the output from the previous quanted block or the previous original block.
@@ 4.3 Comparing With Rounding Methods
In this section, we conduct a comparative analysis between SignRound, FlexRound(Lee et al., 2023), and AdaRound(Nagel et al., 2020). Notably, there is no formal official implementation available for FlexRound and AdaRound for LLMs. Hence, we reference the code  [^3] https://openreview.net/forum?id=-tYCaP0phY_ for further details. However, it’s important to highlight that due to the lack of AMP support and other optimizations, the implementation is notably slow, especially when adhering to the official settings, which involve tuning 5000 steps, as presented in Table 9. Hence, our comparison is limited to models of size 13B or smaller. We set the learning rate to 2e-4 for LLaMA-2-7b and Mistral-7B, and 1e-4 for LLaMA-13-7b to align with the official settings as closely as possible.
@@
Model Method Iter. Mistral-7B V2-7B V2-13B
@@ 4.4 Ablation Study
To validate the effectiveness of SignSGD, Table 4 compares it with the Adam optimizer. SignSGD employs a fixed learning rate of $35e-$ 3 throughout all experiments, comprising 200 steps, with linear weight decay. For Adam, we explored initial learning rates ranging from $32.5e-$ 3 to $22e-$ 2. We choose to quantize models of 13B or less with W4G-1 due to the experiment’s cost. SignSGD demonstrated a distinct advantage in average accuracy metrics across 11 tasks, which demonstrate the unique advantage of signed gradient descent in this scenario.
@@
Round and Weigh clip tuning
To validate the contributions of rounding tuning and weight clip tuning, we conducted ablation studies on three models with two quantization configurations. As shown in Table 5, each component provides benefits over RTN, with rounding tuning offering greater advantages. However, when combined, weight clip tuning can sometimes result in lower accuracy in some cases at W2G128.
@@
Hypeparameters sensitivity
To validate the sensitivity of hyperparameters in SignRound, we conducted ablation studies on sequence length for calibration, number of samples for calibration, tuning batch size, tuning steps, and tuning learning rate. The results are presented in Table 6. Overall, our default hyperparameters achieved balanced results.
@@
Model Seqlen_512 Samples_128 Batch_4 Iters_100 LR_1e-2 Default
Mistral-7B 60.32 61.82 61.78 61.06 61.27 62.33
V2-7B 57.91 56.41 57.21 57.10 55.89 57.48
V2-13B 60.88 60.87 61.21 60.80 61.03 61.20
@@ 4.5 Generalization to other models
To assess the generalization of our method on LLMs, we evaluate SignRound on various mainstream LLMs like Gemma (Team et al., 2024), Phi (Li et al., 2023b), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024) and Llama3 (Touvron et al., 2024). Table 7 demonstrated that all int4 models met the requirement of maintaining accuracy drop within 1% of the FP16 or BF16 accuracy by employing 1000 tuning steps and model wise hype-parameters among 4 choices detailed in scection 4.1. Additionally, one-third of the models exhibited a slight improvement in accuracy compared to the baseline. Please note that the generalization experiments utilize an updated version (0.4.0+) of lm-eval-harness (Gao et al., 2023) and real quantized model, which may result in minor discrepancies compared to other benchmark data.
@@
5 Conclusions and Limitations.
In this paper, we introduce SignRound, an efficient and succinct approach to optimizing weight rounding in quantization tasks. SignRound leverages signed gradient descent for lightweight block-wise tuning, resulting in significant quantization enhancements with minimal iteration steps. Our extensive experiments demonstrate that SignRound outperforms other quantization methods across various models and evaluation metrics while maintaining fixed hyperparameter configurations and similar fine-tuning time costs. Additionally, SignRound offers increased flexibility in quantizing specific large language models, achieving enhanced quantization performance with appropriately tuned hyperparameters.
Despite these significant advantages, we still observed a noticeable gap in the accuracy performance of ultra-low bit quantization compared to the original model, such as 2-bit quantization. This challenge could potentially be addressed by exploring non-uniform quantization and mixed-precision quantization, which we leave it as the future work.
Appendix A Quantiztion Cost
Table 8 compares the quantization costs of different methods, with all measurements conducted on a single NVIDIA A100 GPU with 80GB of memory. We ensure each evaluation process exclusively occupies one GPU, but CPU and other resources may be shared among different processes due to limited resource. For SignRound, we disabled low_gpu_mem_usage in our implementation to achieve faster tuning, albeit with higher memory usage. Nevertheless, LLaMAV2-70B was still able to run on an A100 GPU with 80GB of memory. Although HQQ is exceptionally fast, our methods outperform others in terms of speed. Table 9 also compares the costs between FlexRound, Adaptive Round, and our method.
@@ Appendix A Quantiztion Cost
Method FlexRound AdativeRound Ours
@@ Appendix B View of distribution of tuned parameters
Table 14: Perplexity(PPL) ($\downarrow$) of Wikitext2, PTB , C4 and Wikitext tasks for LLaMA and Mistral models. we follow the source code of GPTQ for wikitext2, PTB and C4 PPL evaluation, while for wikitext, we adopt lm-eval-harness [Gao et al., 2023]. NAN indicates no data acquisition, while OOM denotes out of memory.
@@ Appendix C Other results
We present the detailed 11 tasks accuracy results for llama and mistral models with a size of 7-70B at W2-W4 in Table 10, 11, 12 and 13. the detailed PPL results in Table 14. Overall, our SignRound demonstrates a clear advantage in accuracy tasks, particularly in ultra-low bit quantization, achieving state-of-the-art performance compared to several popular weight quantization methods. In terms of perplexity (PPL), SignRound outperformed all other methods in 57 out of 112 scenarios, demonstrating its advantages. However, we observed that several quantization algorithms, including SignRound, exhibit sensitivity across different models and tasks. The reason for this sensitivity is detailed in Section 4.1.

### v4

@@ Abstract
Large Language Models (LLMs) have demonstrated exceptional proficiency in language-related tasks, but their deployment poses significant challenges due to substantial memory and storage requirements. Weight-only quantization has emerged as a promising solution to address these challenges. Previous research suggests that fine-tuning through up and down rounding can enhance performance. In this study, we introduce SignRound, a method that utilizes signed gradient descent (SignSGD) to optimize rounding values and weight clipping within just 200 steps. SignRound integrates the advantages of Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ), achieving exceptional results across 2 to 4 bits while maintaining low tuning costs and avoiding additional inference overhead. For example, SignRound achieves absolute average accuracy improvements ranging from 6.91% to 33.22% at 2 bits. It also demonstrates robust generalization to recent models and achieves near-lossless quantization in most scenarios at 4 bits. The source code is publicly available at https://github.com/intel/auto-round.
@@ 4 Experiments
This section presents a comprehensive evaluation of SignRound from various perspectives. We begin with a concise overview of the LLM architectures and tasks included in our assessment. Next, we provide a detailed comparison between our method and several existing approaches, emphasizing the unique advantages of SignRound. Furthermore, we conduct ablation studies to reinforce the efficacy of our choices and investigate the sensitivity of hyperparameters. Lastly, we evaluate the generation ability of our method across various recent models. The tuning cost comparisons are provided in Appendix A.
@@ 4.1 Experimental Settings
We evaluate multiple language tasks to address the task-agnostic setting. Specifically, we present the average accuracy results for 11 zero-shot tasks, including HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), LAMBADA (Paperno et al., 2016), TruthfulQA (Lin et al., 2021), OpenBookQA (Mihaylov et al., 2018), BoolQ (Clark et al., 2019), RTE (Dagan et al., 2010), ARC-Easy, ARC-Challenge (Clark et al., 2018), and MMLU (Hendrycks et al., 2020). We use lm-eval-harness (Gao et al., 2023) for all the above tasks. Furthermore, we complement our evaluation with perplexity (PPL) analysis on Wikitext2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020), following the source code [^1] https://github.com/IST-DASLab/gptq of GPTQ and Wikitext2 (Merity et al., 2016) using lm-eval-harness (Gao et al., 2023). However, we argue that perplexity is notably influenced by outliers, as illustrated in Table 14 for different algorithms. This susceptibility likely arises from the mathematical expression $\mathrm{PPL}(X)=\exp\left(-\frac{1}{t}\sum_{i=1}^{t}\log p_{\theta}(x_{i}|x_{<i})\right)$, where assigning a low probability to even one token can significantly inflate the perplexity score. Consequently, we prioritize the accuracy of the 11 tasks mentioned above as the primary metric, with perplexity data serving as supplementary reference.
@@
In alignment with GPTQ (Frantar et al., 2022), our focus is specifically on weight-only quantization, targeting the linear layers within transformer blocks. Layers such as the embedding layer and typically the last linear layer like ’lm-head’ are excluded from the quantization process. Our evaluation primarily centers on W4, W4G128, and W3G128 configurations, where W4 indicates quantizing weights with 4 bits and G represents finer-grained grouping as described in (Park et al., 2022; Frantar et al., 2022). We adopt asymmetric quantization. To mitigate overfitting on the WikiText and C4 datasets, we randomly select 512 samples with the same seed from the readily available pile-10k dataset [^2] https://huggingface.co/datasets/NeelNanda/pile-10k for calibration, which comprises the first 10k samples from pile (Gao et al., 2020). We used a sequence length of 2048 for calibration, while for other methods, we adhere to their official settings.
@@
We compare different algorithms on commonly used models such as LLaMA-V1 (Touvron et al., 2023a), LLaMA-V2 (Touvron et al., 2023b), and Mistral-7B-v0.1 (Jiang et al., 2023). Our comparison covers a wide range of LLM parameters, ranging from 7B to 70B, to ensure comprehensive coverage and analysis.
Table 1: Average accuracies ($\uparrow$) across 11 tasks, as detailed in Section 4.1, for LLaMA and Mistral models at W2-W4. Our* denotes the highest accuracy achieved among the 8 hyperparameter choices, outlined in Section 4.2, whereas for the 70B model, we tested only a few options.
@@
Table 2: Detailed accuracies($\uparrow$) across 11 tasks(0-shot) of LLaMA and Mistral models at W4G-1. Our* denotes the highest accuracy achieved among the 8 hyperparameter choices, outlined in Section 4.2, whereas for the 70B model, we tested only a few options. Appendix C provides more detailed data.
@@
SignRound Hyperparameters.
Unless explicitly stated, the tuning process involved adjusting each block for 200 steps with a learning rate of $1/200$, i.e., $35e-$ 3, a batch size of 8, and linear learning rate decay. Additionally, we employed automatic mixed precision (AMP) to accelerate the tuning.
@@ 4.2 Comparing With Recent Methods
In this section, we compare our methods with those that have already demonstrated remarkable results and impose no additional overhead on our tested models in weight-only quantization for LLMs, including GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2023), HQQ (Badri and Shaji, 2023), OmniQuant (Shao et al., 2023) with a naive method RTN.
To ensure fair comparison as much as possible, we enabled act-order and true-sequential modes in GPTQ and also activated static_group in scenarios with group_size. The notation GPTQ+ indicates that we adjusted the random seed or data pre-processing to address issues related to the non-positive definite Hessian matrix or other issues. For OmniQuant(Shao et al., 2023), we adhere to the official settings, which include running for 20 epochs including W2G128 for saving time and disabling ’let’. We conducted calibration tests using sample sizes of 512 and 128, as well as a sample size of 512 with a batch size of 4. Our findings show that using a sample size of 512 typically results in comparable or slight higher performance for models less than or equal to 13B. Therefore, we present the results based on the sample size of 512. For 70B models, due the the NAN loss issue and to reduce the tuning cost of OmniQuant, we adopted 128 samples for calibration.
We present the summary results of Mistral-7B and LLAMAV2 in Table 1 , detailed results of W4G-1 in Table 2, and additional detailed results are provided in Appendix C due to space constraints. In summary, our approach demonstrated superior performance compared to GPTQ (Frantar et al., 2022), achieving scores of 30/32, AWQ (Lin et al., 2023) with 27/32, HQQ (Badri and Shaji, 2023) with 15/16, and OmniQuant (Shao et al., 2023) with a score of 29/32 across llamv1/llamav2/mistral-7b on various quantization settings, including W4G-1, W4G128, W3G128, and W2G128. These evaluations were based on the average accuracies of 11 zero-shot tasks.
It’s worth noting that as the bit depth decreases, the advantages of SignRound become more notable. For example, at W2G128, SignRound could yield absolute average accuracy improvements ranging from 6.91% to 33.22%.
Moreover, we can enhance the performance by optimizing the model’s hyperparameters from a selection of eight choices, denoted as ours*. These choices include steps (200, 1000), weight clip learning rate (1.0/steps, 2.0/steps), and the option to either enable or disable quantized inputs, which refers to utilizing the output from the previous quantized block or the previous original block.
@@ 4.3 Comparing with Rounding Methods
In this section, we conduct a comparative analysis between SignRound, FlexRound(Lee et al., 2023), and AdaRound(Nagel et al., 2020). Notably, there is no formal official implementation available for FlexRound and AdaRound for LLMs. Hence, we reference the code  [^3] https://openreview.net/forum?id=-tYCaP0phY_ for further details. However, it’s important to highlight that due to the lack of AMP support and other optimizations, the implementation is notably slow, especially when adhering to the official settings, which involve tuning 5000 steps, as presented in Table 9. Therefore, our comparison is limited to models of size 13B or smaller. We set the learning rate to 2e-4 for LLaMA-2-7b and Mistral-7B, and 1e-4 for LLaMA-13-7b to align with the official settings as closely as possible.
@@
Model Method Steps Mistral-7B V2-7B V2-13B
@@ 4.4 Ablation Studies
To validate the effectiveness of SignSGD, Table 4 compares it with the Adam optimizer Kingma and Ba (2014). SignSGD employs a fixed learning rate of $35e-$ 3 throughout all experiments, comprising 200 steps, with linear weight decay. For Adam, we explored learning rates ranging from $32.5e-$ 3 to $22e-$ 2. We choose to quantize models of 13B or less with W4G-1 due to the experiment’s cost. SignSGD demonstrated a distinct advantage in average accuracy metrics across 11 tasks, which demonstrate the unique advantage of signed gradient descent in this scenario.
@@
Round and Weigh Clip Tuning.
To validate the contributions of rounding tuning and weight clip tuning, we conducted ablation studies on three models with two quantization configurations. As shown in Table 5, each component provides benefits over RTN, with rounding tuning offering greater advantages. However, when combined, weight clip tuning can sometimes result in lower accuracy in certain cases at W2G128.
@@
Hypeparameters Sensitivity.
To validate the sensitivity of hyperparameters in SignRound, we conducted ablation studies on sequence length for calibration, the number of samples for calibration, tuning batch size, tuning steps, and tuning learning rate. The results are presented in Table 6. Overall, our default hyperparameters achieved balanced results.
@@
Model Seqlen_512 Samples_128 Batch_4 Steps_100 Steps_1000 LR_1e-2 Default
Mistral-7B 60.32 61.82 61.78 61.06 62.58 61.27 62.33
V2-7B 57.91 56.41 57.21 57.10 57.19 55.89 57.48
V2-13B 60.88 60.87 61.21 60.80 61.01 61.03 61.20
@@ 4.5 Generalization to Other Models
To assess the generalization of our method on LLMs, we evaluate SignRound on various mainstream LLMs such as Gemma (Team et al., 2024), Phi (Li et al., 2023b), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024) and Llama3 (Touvron et al., 2024). Table 7 demonstrated that all int4 models met the requirement of maintaining accuracy drop within 1% of the FP16 or BF16 accuracy by employing 1000 tuning steps and model wise hype-parameters among 4 choices detailed in section 4.1. Additionally, one-third of the models exhibited a slight improvement in accuracy compared to the baseline. Please note that the generalization experiments utilize an updated version (0.4.0+) of lm-eval-harness (Gao et al., 2023) and real quantized model, which may result in minor discrepancies compared to other benchmark data.
@@
5 Conclusions and Limitations
In this paper, we introduce SignRound, an efficient and concise approach for optimizing weight rounding in the quantization of large language models. SignRound employs signed gradient descent for tuning rounding value and weight clipping in 200 steps, completing the quantization of LLAMA-V2-70B in approximately 2.5 hours. Our extensive experiments show that SignRound outperforms other quantization methods across various models and weight bits, maintaining fixed hyperparameter settings. Additionally, SignRound shows promising generation capabilities in recent models and achieves enhanced performance through model-specific hyperparameter tuning.
Despite these significant advantages, we still observed a noticeable gap in the accuracy performance of ultra-low bit quantization compared to the original model, such as 2-bit quantization. This challenge could potentially be addressed by exploring non-uniform quantization and mixed-precision quantization, which we leave for future work.
Appendix A Quantization Cost
Table 8 compares the quantization costs of different methods, with all measurements conducted on a single NVIDIA A100 GPU with 80GB of memory. We ensure each evaluation process exclusively occupies one GPU, but CPU and other resources may be shared among different processes due to limited resources. For SignRound, we disabled low_gpu_mem_usage in our implementation to achieve faster tuning, albeit with higher memory usage. Despite this, LLaMAV2-70B was still able to run on an A100 GPU with 80GB of memory. Although HQQ is exceptionally fast, our methods outperform others in terms of speed. Table 9 also compares the costs between FlexRound, Adaptive Round, and our method.
@@ Appendix A Quantization Cost
Method FlexRound AdaRound Ours
@@ Appendix B View of distribution of tuned parameters
Table 14: Perplexity(PPL) ($\downarrow$) of Wikitext2, PTB , C4 and Wikitext tasks for LLaMA and Mistral models. we follow the source code of GPTQ for wikitext2, PTB and C4 PPL evaluation, while for wikitext, we adopt lm-eval-harness [Gao et al., 2023]. NAN indicates not a number, while OOM denotes out of memory.
@@ Appendix C Other results
We present the detailed accuracy results for 11 tasks using the LLaMA and Mistral models, ranging in size from 7B to 70B, at W2-W4 in Tables 10, 11, 12 and 13. The detailed perplexity (PPL) results are shown in Table 14. Overall, SignRound demonstrates a clear advantage in accuracy tasks, particularly in ultra-low bit quantization, achieving state-of-the-art performance compared to several popular weight quantization methods. In terms of perplexity (PPL), SignRound outperformed all other methods in 57 out of 112 scenarios, demonstrating its advantages. However, we observed that several quantization algorithms, including SignRound, exhibit sensitivity across different models and tasks. The reason for this sensitivity is detailed in Section 4.1.